Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #6725

Period Covered:
07:00 pm December 11, 2014 - 06:59 pm December 12, 2014
(DOY 346/0000z - 346/2359z)

Flight Operations Summary

This report was generated by the HST Automated Report System @2014-12-13 17:02:00 GMT

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)

HSTARs:
none

COMPLETED OPS REQUEST:
none

COMPLETED OPS NOTES:
none

FGS ACQ STATUS:
		    SCHEDULED       SUCCESSFUL
FGS GSAcq		6		6
FGS REAcq		10		10
OBAD with Maneuver	6		6

SIGNIFICANT EVENTS:
At 345/13:14, Ops Request 19714 was started to update the CPM3 EEPROM via Ops Request 19714.  
Unfortunately, a hardware parity error was received at 345/13:21at the start of the copy from CPM1 
to CPM3 Bank 0.   
 
As discussed in the FRR contingency planning, the copy was aborted, a hardware reset was sent, and 
the copy was attempted again.  The hardware parity error was again received, but the team continued 
with the procedure to complete the copy and collect a memory dump of CPM3 Bank 0.  Additional 
parity errors were received during attempts to get memory dumps from CPM3 Bank 0 and Bank 1.   
 
The team decided to stand down from further investigation activities today, and CPM3 was powered 
off per the nominal procedures at 19:21.   
 
The NSSC-1 FSW version 9.2.3 image resident in the 486 was successfully loaded, dumped, and 
verified.  The appropriate QBF loads have been sent to the spacecraft and QBF was re-enabled.  
Nominal automated operations have resumed. 
 
HSTAR 14207 was written to document the anomaly.  A Tiger Team will be formed to investigate the 
CPM3 errors. 



Program Principal Investigator Program Title
Science Observations Scheduled
13382 Mary Putman, Columbia University in the City of New York Warm Gas Flows in the Coma Cluster
13459 Tommaso Treu, University of California - Los Angeles The Grism Lens-Amplified Survey from Space {GLASS}
13504 Jennifer Lotz, Space Telescope Science Institute HST Frontier Fields - Observations of MACSJ1149.5+2223
13644 Xiaohui Fan, University of Arizona CIII] Emission in z=5.7 Galaxies: A Pathfinder for Galaxy Spectroscopy in the Reionization Era
13743 David Thilker, The Johns Hopkins University The Controversial Nature of the Diffuse UV Emission in Galaxies: Exploring NGC300
13761 Stephan McCandliss, The Johns Hopkins University High efficiency SNAP survey for Lyman alpha emitters at low redshift
13793 Rebecca Bowler, Royal Observatory Edinburgh Unveiling the merger fraction, sizes and morphologies of the brightest z ~ 7 galaxies
Calibration Observations Scheduled
13602 Ralph Bohlin, Space Telescope Science Institute Bandpass variations across the FOV
13952 David Golimowski, Space Telescope Science Institute CCD Daily Monitor (Part 1)
13968 Justin Ely, Space Telescope Science Institute FUV Detector Dark Monitor
13974 Justin Ely, Space Telescope Science Institute NUV Detector Dark Monitor
13980 Joanna Taylor, Space Telescope Science Institute CCD Dark Monitor Part 1
13982 Joanna Taylor, Space Telescope Science Institute CCD Bias and Read Noise Monitor Part 1
14002 Matthew Bourque, Space Telescope Science Institute WFC3 UVIS Daily Monitor A
14030 Peter McCullough, Space Telescope Science Institute WFC3 CSM monitor Earth Flats